sum of interior angles of a polygon with 5 sides
sum of interior angles of a polygon with 10 sides
sum of interior angles of a polygon with 4 sides
sum of interior angles of a polygon with 7 sides

Answer by Nate(3500)

The sum is 180(n - 2) for any n-sided polygon
5 sides means n = 5
180(5 - 2) = 180 * 3 = 540 degrees